OpenStudy (anonymous):

Help please. Write an equation that describes each line in slope intercept form. Slope 4 (-3,4) is on the line.

hartnn (hartnn):

use this \(\Large y-y_1 = m (x-x_1 )\) here , m = 4 x1 = -3 y1= 4

OpenStudy (anonymous):

Ok, thank you.

hartnn (hartnn):

welcome ^_^